# 2TRIUMF, 4004 Wesbrook Mall, Vancouver, BC V6T 2A3 Canada

### v1

@@ V Conclusions
We have used our $\beta$ asymmetry data from the decay of 37K to constrain new physics that depends on the $\beta$ energy, measuring the Fierz interference term from quark-lepton Lorentz scalar and vector couplings $b_{F}$ = -0.0002 $\pm$ 0.024 (stat) $\pm$ 0.039 (syst), We provide constraints on a combination of Lorentz scalar and tensor lepton-quark couplings. Our present results using our laser-cooled atoms are dominated by a combination of statistics and solvable $\beta$ scattering, and we hope to improve sensitivity by a factor of about five to enable complementarity with experiments indicating possible existence of non-standard model physics.
@@ Acknowledgements.
We acknowledge TRIUMF/ISAC staff, in particular for TiC target preparation. Supported by the Natural Sciences and Engineering Research Council of Canada, and the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Awards No. DE-FG03- 93ER40773 and No. DE-FG02-11ER41747. TRIUMF receives federal funding via a contribution agreement through the National Research Council of Canada.

### v4

@@ V Conclusions
We have used our $\beta$ asymmetry data from the decay of 37K to constrain new physics that depends on the $\beta$ energy, measuring the Fierz interference term $b$ = 0.033 $\pm$ 0.084 (stat) $\pm$ 0.039 (syst). We also provide constraints on an exclusion plot on a combination of Lorentz scalar and tensor lepton-quark couplings from our determination of $b$ and $A_{\beta}$ adding the measured $fT$ of 37K and its uncertainty [19] (using the center value of $fT$ of $0^{+}\rightarrow 0^{+}$ decays for the Fermi strength). Our present results using our laser-cooled atoms are dominated by a combination of statistics and solvable $\beta$ scattering, and we hope to improve our statistics-limited result by an order of magnitude to enable complementarity with experiments indicating possible existence of non-standard model physics.
@@ Acknowledgements.
We acknowledge physics input from Martín González-Alonso. We acknowledge TRIUMF/ISAC staff, in particular for TiC target preparation. Supported by the Natural Sciences and Engineering Research Council of Canada, and the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Awards No. DE-FG03- 93ER40773 and No. DE-FG02-11ER41747. TRIUMF receives federal funding via a contribution agreement through the National Research Council of Canada.